OpenStudy (anonymous):

The population P of a city if given by P=2500e^(kt) where t is time of the years , with t=0 corresponding to the years 1980 .In 1935 ,the population was 3350 .Find the value of k and use this result to predict the population in the years 2000

OpenStudy (anonymous):

I'm sorry are you sure that isn't 1985? instead of 1935??? You are beginning time with 1980 (t=0)...I'm not sure it would make sense to have t=-45 in the year 1935. I think you have 1985. If so then you have to plug these values into the equation: 3350=2500e^k(5) (In 1980 time t=0 so in 1985 5 years have passed by and t is in years) Then we have 3350/2500=e^5k ln(3350/2500)=5k ln(3350/2500)/5=k Now plug in k in your given equation P=2500e^(kt). Remember you are looking for the population in the year 2000 so your t is now t=2000-1980.

OpenStudy (anonymous):

yes,I positive I think is decay because is -45 years